# A Theory of Quantum Jumps and Its Applications to Lyman and Balmer Series

### v3

@@ Abstract
A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation. The new equation displays Bohr’s rules for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent. The temporal solutions are used to determine time-scales of the quantum jumps, which are very short but finite. The spatial solutions are used to find the radial probability density of finding the electron after its jump. The obtained results are applied to the Lyman and Balmer series absorption, and their experimental verification is discussed.
@@ 1 Introduction
The randomness and abruptness of quantum jumps between two discrete energy levels in an atom was the main feature of quantum theory originally formulated by Bohr [1] in 1913. His main argument was that electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, and that the quantum jumps corresponding to absorption and emission are instantaneous. This notion of random and instantaneous quantum jumps became one of the pillars foundation of the Copenhagen interpretation of quantum mechanics (QM) (e.g., [2-5]). However, Schrödinger [6] argued against such instantaneous quantum jumps by pointing out that no process in the real world happens instantenously or in zero time.
@@
The first direct observations of quantum jumps in individual atoms exposed to electromagnetic (EM) fields took place in 1986 and were reported in [7-9]. The collected data showed that the atoms flipped between a state that allowed them to emit a photon to a state that did not, and that the atoms remain in one state, or the other, for periods ranging from a few tenths of a second to a few seconds before jumping again. Other indepedent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein). A different view has emerged from the experimental results obtained by Minev et al [13], who confirmed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that unfolds in a finite time. It was suggested that a quantum trajectories theory could account for the experimental data [14-16].
@@
The theoretical approach described in this paper is different from any previous attempt to account for quantum jumps, including the quantum trajectories theory. The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) that was derived from the irreducible representations (irreps) of the extended Galilean group [17]. Since the same irreps can be used to derive the Schrödinger equation (SE), and since there are other similarities between the equations, the SE and NAE are considered to be complementary to each other. The main physical implication of this complementarity is that quantum phenomena, like quantum jumps or quantum measurements, are not described by the SE, but they may be represented by the NAE.
@@
The NAE naturally describes interactions of quantum systems with environments that are subjected to monitoring by measuring devices. Specifically, the NAE can directly describe interactions of a single quantum particle with a measuring apparatus, or the process of absorption or emission of electromagnetic (EM) waves by quantum systems. Thus, while the SE is used to formulate unitary quantum theories, the NAE can also be used to formulate non-unitary theories because of its second-order derivatives in time (see Section 2, for details). Random and instantaneous quantum jumps violate the unitary evolution (e.g., [2-13]), thus, an attempt herein to describe them by using the NAE. Similarly, the quantum measurement problem is a non-unitary evolution of a quantum system during its measurement (e.g., [18-21]) and, therefore, the NAE was used to solve it [22].
It was recently demonstrated that a non-unitary theory of quantum measurements can be formulated using the NAE, and that its predictions are consistent with Born’s position measurement postulates [22]. Moreover, since the origin and nature of dark matter remains unknown (e.g., [23-26]), and since all previous attempts to formulate a theory of dark matter (DM) based on the SE (e.g., [27,28]) failed (e.g., 29,30]), a theory of DM using the NAE was developed [22], and applied to construct models of galactic cold DM halos [31,32], whose theoretical predictions are proposed to be verifed by specific astronomical observations [33]. The NAE was also used to describe propagation of classical (acoustic) waves in isothermal and stratified media [34].
@@
The paper is organized as follows: the basic equations resulting from the Galilean group are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 7.
@@ 2 Schrödinger and new asymmetric equations
In Galilean relativity, space and time are separated and described by different metrics, which are invariant with respect to rotations, translations and boosts that form the Galilean group (e.g., [35,36]). The irreducible representations (irreps) of the Galilean group were originally obtained and classified by Bergmann [37], who demonstrated that only scalar and spinor irreps are physical in Galilean space and time. The spinor irreps were used by Lévy-Leblond [38,39] to obtain his spinor wave equation that describes elementary particles with spins in nonrelativistic QM. In this paper, only the scalar irreps are considered.
In the extended Galilean group (e.g., [38,39]), the space and time translations are combined into one subgroup, $T(3+1)$, whose properties are the same as the $T(3+1)$ subgroup of the Poincaré group [35,36]. Since $T(3+1)$ is an invariant subgroup, its irreps are the irreps of the entire extended Galilean group, and they can be used to derive equations of nonrelativistic and relativistic QM [40]. According to Wigner [41], an elementary particle is uniquely identified if its wavefunction transforms as one of the irreps of the extended Galilean group. For a scalar wavefunction $\phi(t,\mathbf{x})$, the Wigner condition is given by the following eigenvalue equations $i\partial_{t}\phi=\omega\phi$ and $-i\nabla\phi=\mathbf{k}\phi$, where $\partial_{t}=\partial/\partial t$, and $\omega$ and $\mathbf{k}$ are the eigenvalues that label the irreps [42,43].
Using the eigenvalue equations and the de Broglie relationship (e.g., [3-5]), the following two asymmetric equations are derived [17,32]:
@@
which is the Schrödinger equation (e.g., [3-5]), and
@@
which is called the new asymmetric equation or NAE [31-33]. Since both equations are derived from the same irreps, they may be considered as complementary to each other. However, the main difference between them is that the SE is the first-order in its time derivative, thus allows for unitary theories, while the NAE can also be used to construct non-unitary theories because of its second-order time derivative. In addition, the NAE requires specifying the eigenvalues $\omega$ and $\mathbf{k}$ that are explicitly present in the equation.
There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps, that are considered to be non-unitary [3-5]. In these processes, the external input/output of EM radiation plays important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$. A possible solution of the quantum measurement problem by using the NAE is presented in [22]. In this paper, the NAE is used to formulate a theory of quantum jumps.
3 Basic equations and their solutions
@@ 3.1 The governing equation
where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential.
To determine $\omega_{o}$ and ${\mathbf{k_{o}}}$, one may use
@@
where $R_{E}=\hbar^{2}/2ma_{B}^{2}$ is the Rydberg constant, $a_{B}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}$ is the Bohr radius, and $n_{i}$ and $n_{f}$ are the prinicipal quantum numbers corresponding to the initial and final states, respectively. For absorption, $n_{f}>n_{i}$ and $\Delta E>0$; however, for emission, $n_{f}<n_{i}$ and $\Delta E<0$. Taking $\Delta E=\hbar\omega_{o}$, one obtains
$\omega_{o}=\frac{R_{E}}{\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (5)
@@
${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{c\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right){\mathbf{\hat{k}}_{o}}\ .$ (6)
Then, Eq. (3) can be written as
@@ 3.2 Temporal and spatial solutions
To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $-\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$. Then, Eq. (7) becomes
@@
where $n$ is the principle quantum number, and $n=n_{f}$ for absorption and $n=n_{i}$ for emission; since the solutions also depend on $n$, one gets $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$.
@@
${\cal{R}\it e}[\chi_{n}(t)]=C_{+}\cos{(\Omega_{n}t)}\exp{(-\Omega_{n}t)}\ ,$ (12)
and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 3.
Before solving Eq. (9), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$. It is also convenient to transform Eq. (9) to the new independent variable $r_{a}=r/a_{B}$, which gives
@@
According to Eq. (13), $\eta_{n}(r)$ reaches its maximum value when $r_{a}=n^{2}$ or $r=n^{2}a_{B}$, which is Bohr’s rule for quantum jumps (e.g., [3-5]). In other words, Eq. (13) directly displays the Bohr rule that gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
@@
$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{o}}\ \exp{\left[-\left(\frac{\beta_{o}}{n^{2}}\right)r_{a}\right]}\ ,$ (15)
@@
which is valid if ${\cal{R}\it e}(2\beta_{o})>0$ and ${\cal{R}\it e}(4\beta_{o}+2)>-1$; note that both conditions are satisfied in the theory presented herein.

### v4

@@ Abstract
A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation. The new equation is a mathematical form of Bohr’s rule for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent. The temporal solutions are used to determine time-scales of the quantum jumps, which are found to be not instantaneous but finite. The spatial solutions are used to obtain the radial probability density for an electron after its jump. The obtained results are applied to the Lyman and Balmer series, and their experimental verification is discussed.
Abstract
A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation. The new equation is a mathematical form of Bohr’s rule for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent. The temporal solutions are used to determine time-scales of the quantum jumps, which are found to be not instantaneous but finite. The spatial solutions are used to obtain the radial probability density of an electron after its jump. The obtained results are applied to the Lyman and Balmer series, and their experimental verification is discussed.
keywords:
Quantum mechanics, quantum jumps, Lyman and Balmer series
@@ 1 Introduction
Quantum jumps are transitions of electrons between discrete energy levels in atoms when electromagnetic radiation of a specific energy is absorbed or emitted by the atoms. In the original hydrogen atom model [1], Bohr postulated that the passing between different quantum states in the atom, when electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, cannot be properly described by mechanics. Einstein [2] introduced the notion that such transtions are random and questioned their randomness, while Bohr responded by arguing that the transitions are instantaneous. As pointed out by Schrödinger [3], later the transitionswere named ’quantum jumps’ or were also called Bohr-Einstein quantum jumps. Schródinger [3] argued against the jumps being instantaneous as no process in the real world happens instantenously or in zero time. Then, the idea of quantum jumps being random and instantaneous, which has been reapetedly verified experimentally, has become associated with the Copenhagen interpretation of quantum mechanics (e.g., [2-5]).
@@
The first direct observations of quantum jumps in individual atoms exposed to electromagnetic (EM) fields took place in 1986 and were reported in [7-9]. The collected data showed that the atoms flipped between a state that allowed them to emit a photon to a state that did not, and that the atoms remain in one state, or the other, for periods ranging from a few tenths of a second to a few seconds before jumping again. Other independent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein). A different view has emerged from the experimental results obtained by Minev et al. [13] confirming that quantum jumps occur at random times, but once they occur, the evolution of each completed jump appears to be a continuous and coherent physical process that unfolds in a finite time.
There were many attempts to formulate theories of quantum jumps in atomic systems aiming to relate the resulting theoretical predictions to the experimental data available at the time when such theories were developed (e.g., [14-19]). Specifically, it was suggested that the so-called quantum trajectory theory (QTT) may account for the experimental data presented in [13]. QTT gives trajectories of individual particles that obey the probabilities computed from the Schrödinger equation [20,21], and the theory is used to study both open and dissipative quantum systems. There are three different quantum trajectories considered in QTT: non-real quantum trajectories treated as numerical tools [20], and subjectively [21] and objectively [22] real quantum trajectories (e.g., [23]). A trajectory in QTT is not like a classical trajectory, but instead it is a path in Hilbert space that is determined by solving the stochastic Schrödinger equation (e.g., [23]). Predictions made by QTT cannot be reproduced by the standard quantum mechanics approach based on the Schrödinger equation (e.g., [24]).
More recently, a theory of quantum jumps was developed for atoms that are coupled to a quantized electromagnetic field in the limiting regime where the orbital motion of the atoms is neglected and the velocity of light is assumed to be infinitely large [25]. It must also be pointed out that quantum jumps play important roles in quantum feedback control [26,27], and in detecting and fixing decoherence-induced errors in quantum information systems [28,29].
@@
The theoretical approach presented in this paper is very different from any previous attempt to account for quantum jumps. The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) [30] that was derived from the irreducible representations (irreps) of the Galilean group of the metrics (e.g., [31-33]). Since the same irreps were used to derive the Schrödinger equation [34,35], the equations are considered to be complementary to each other [30]. The main physical implication of this complementarity is that some quantum phenomena, like the quantum jumps considered in this paper, or the quantum measurement problem [36], cannot be accounted for by the Schrödinger equation and they require the NAE to describe them.
@@
The NAE naturally describes interactions of quantum systems with environments that are subjected to monitoring by measuring devices. Specifically, the NAE can be used to describe the dynamics of quantum particles during their transitions in atoms as shown in this paper, or interactions of a single quantum particle with a measuring apparatus, as demonstrated in [36], or other physical processes [30]. The main reason is that non-unitary theories of these quantum processes can be developed using the NAE because of its second-order derivatives in time (see Section 2, for details). Random and instantaneous quantum jumps violate the unitary evolution, and they cannot be represented by the Schrödinger equation (e.g., [2-13]); thus, in this paper, a theory of quantum jumps is developed based on the NAE.
@@
The paper is organized as follows: the basic equations resulting from the Galilean group of the metrics are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 6.
@@ 2 Schrödinger and new asymmetric equations
In Galilean relativity, space and time are separated and their metrics are given by $ds_{1}^{2}\ =\ dx^{2}+dy^{2}+dz^{2}$ and $ds_{2}^{2}\ =\ dt^{2}$, where $x$, $y$ and $z$ are spatial coordinates and $t$ is time. The transformations that make the metrics invariant form the Galilean group of the metric, $\mathcal{G}$}, whose structure is ${\mathcal{G}}\ =S(4)\otimes_{s}H(6)$, where $S(4)=T(1)\otimes R(3)$ is a four-parameter subgroup of translation in time and rotations in space, and $H(6)=T(3)\otimes B(3)$ a six-parameter subgroup translations in space and boosts [32,33].
The group $\mathcal{G}$} is a ten-parameter Lie group and $H(6)$ is its invariant Abelian subgroup. It is well-known that the irreducible representations (irreps) of $H(6)$ are one-dimensional [29-31], and they provide labels for all the irreps of $\mathcal{G}$}. Classification of the irreps of this group by Bargmann [29] demonstrated that the scalar and spinor irreps are physical, but the vector and tensor irreps are not because they do not allow for the elementary particle localization. Therefore, in nonrelativistic quantum mechanics (NRQM) only scalar and spinor wavefunctions are allowed and their temporal and spatial evolution is described by the Schrödinger [2-5] and Lévy-Leblond [32,33] equations, respectively. In this paper, only the scalar irreps are considered.
Writing separately the transformations for space translations and boosts, one finds: $\hat{T}_{\mathbf{a}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{a},t)\ =e^{i\mathbf{k}\cdot\mathbf{a}}\phi(\mathbf{r},t)$ and $\hat{B}_{\mathbf{v}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{v},t)\ =\ e^{i\mathbf{k}\cdot\mathbf{v}t}\phi(\mathbf{r},t)$, where $\phi(\mathbf{r},t)$ is a scalar wavefunction, $\mathbf{a}$ represents a translation in space, $\mathbf{v}$ is the velocity of Galilean boosts, and the real vector $\mathbf{k}$ labels the one-dimensional irreps of $H(6)$. Expanding $\phi(\mathbf{r}+\mathbf{a},t)$ and $\phi(\mathbf{r}+\mathbf{v}t,t)$ in Taylor series and comparing the results to the above transformations, one obtains the following eigenvalue equation [34-36]
$i\nabla\phi(\mathbf{}r,t)\ =\ \mathbf{k}\phi(\mathbf{}r,t)\ ,$ (1)
which has the same form for the boosts and translations in space, and it is the necessary condition that the wavefunction $\phi(\mathbf{}r,t)$ transforms as one of the irreps of $\mathcal{G}$}.
Now, $S(4)$ is not an invariant subgroup of $\mathcal{G}$}, thus, the above procedure cannot be applied to it; the fact that $T(1)$ is an invariant subgroup of $S(4)$ does not help because $T(1)$ is not the ’little group’ of $\mathcal{G}$} [31,34]. In the solution proposed in [34], a new function, $\phi_{\omega}(\mathbf{r},t)=\eta(\mathbf{r},t)\phi(\mathbf{r},t)$, where $\eta(\mathbf{r},t)$ is a smooth and differentiable function to be determined, and the generator of translation, $\hat{E}=i\partial_{t}=i\partial/\partial t$, were introduced, with $i\partial_{t}\phi_{\omega}(\mathbf{r},t)\ =\ \omega\phi_{\omega}(\mathbf{r},t)$. Then, it was demonstrated [34] that the Galilean Principle of Relativity requires $\eta(\mathbf{r},t)=\eta^{\prime}(\mathbf{r}^{\prime},t^{\prime})=1$, which means that
$i\partial_{t}\phi(\mathbf{r},t)\ =\ \omega\phi(\mathbf{r},t)\ ,$ (2)
and that this eigenvalue equation supplements Eq. (1). The eigenvalue equations given by Eqs (1) and (2) guarantee that the wavefunction $\phi(\mathbf{r},t)$ transforms as one of the irreps of $\mathcal{G}$}, and that these equations can be used to derive dynamical equations that are consistent with Galilean Relativity.
It must be noted that a customary approach of finding a dynamical equation from the Casimir operator of $\mathcal{G}$} cannot be used here because the operator does not connect Galilean space and time. Thus, requires that $\mathcal{G}$} is modified to allow for such connections, which gives the extended Galilean group, whose structure is $\mathcal{G}_{e}\ =[R(1)\otimes B(3)]\otimes_{s}[T(3+1)\otimes U(1)]$, where $U(1)$ is a one-parameter unitary group (e.g., [32,33]); the structure of this group is similar to the Poincaré group (e.g., [37-39]). There are three Casimir operators of $\mathcal{G}_{e}$, but for a scalar wavefunction only one of them gives the Schrödinger equation. However, the problem with this approach is that $\mathcal{G}_{e}$ is not the group of the Galilean metrics, thus, the derivation of dynamical equations using the Casimir operator of $\mathcal{G}_{e}$ may not be consistent with Galilean Relativity.
In the following, the eigenvalue equations, which are consistent with both the irreps of $\mathcal{G}$ and Galilean Relativity, together with the de Broglie relationship (e.g., [3-5]), are used to obtain the following two asymmetric equations [36]:
@@
which is the Schrödinger equation (SE) (e.g., [3-5]), and
@@
which is called the new asymmetric equation (NAE) [30]. Since both equations are derived from the same irreps, they are considered to be complementary to each other. The equations describe free particles, and they must be quantum particles because of the presence of the Planck constant appears as a coefficient in both equations.
The SE is used to compute the probability of finding a quantum system in its possible states before any interaction of the system with its environment, or any measurements are performed on it. The SE is useful for predicting average measurements of large ensembles of quantum objects. The NAE describes and provides insight into the behaviour of individual particles affected by their surroundings represented by $\omega$ and $\mathbf{k}$ in Eq. (4); this treatment of individual particles shows some similarities between QTT and the theory presented in this paper.
Another important difference between the SE and NAE is that the former is first-order in its time derivative, thus allowing for unitary theories, while the latter can also be used to construct non-unitary theories because of its second-order time derivative. In addition, the NAE requires that the eigenvalues $\omega$ and $\mathbf{k}$, which are explicitly present in the equation, are specified; this means that the NAE naturally allows to include quantum system’s interaction with its environment. In other words, by specifying these eigenvalues, an inertial observer becomes an active Galilean observer who changes the system (e.g., performs a measurement) and uses the NAE to describe the resulting changes, which cannot be done by a passive Galilean observer who uses the SE.
There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps that are considered to be non-unitary [3-5]. In these processes, the external input/output of EM radiation plays an important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$. In this paper, the NAE is used to formulate a theory of quantum jumps.
3 Model of quantum jumps
3.1 Formulation
To model quantum jumps, a single hydrogen atom with its electron in the ground state is considered. Quantum jumps are casual because they are triggered by EM radiation, and the NAE given by Eq. (4) is used to describe them. The NAE predicts a new state of the electron after it jumps from the initial energy level, which is one of the s-orbitals, and it gives the duration time for the jump, and also finding the radial probability density of the electron at its new state. Now, for the electron’s final state, it is assumed that the electron returns to its original state either immediately after the jump, or after some time that can be established experimentally.
The considered model of quantum jumps is based on the NAE with the Coulomb potential included in it. Moreover, some results originally obtained from the SE with the Coulomb potential are also used in solving the NAE; specifically, the hydrogen atom energy levels, and the frequencies and wavevectors of EM radiation that are used to trigger quantum jumps. As shown in the previous section, the SE and NAE are complementary, and and both equations contribute to the model. However, the main difference is that the SE describes quantum unitary processes, while the NAE with its second-order time derivative accounts properly for non-unitary processes, such as quantum jumps.
@@ 3.2 The governing equation
where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential, which is included into the NAE to accout properly for the energy levels in the hydrogen atom. Since the SE is used to compute the probability of finding a quantum system in its possible state before any interaction with the system’s environment or any measurement performed on the system, the NAE can be used to describe the system that interacts with its environment by specifying $\omega_{o}$ and $\mathbf{k_{o}}$}; note that if $\omega_{o}=0$ and $\mathbf{k_{o}}=0$}, then the system must be described by the SE. This shows the complementary role these two equations play in NRQM.
The parameters $\omega_{o}$ and $\mathbf{k_{o}}$} in the above equation are determined by using the solutions to the time-independent SE, which provide the wavefunctions and the energy levels associated with these wavefunctions. The energy levels are given by
@@
where $R_{E}=\hbar^{2}/2ma_{B}^{2}$ is the Rydberg constant, $a_{B}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}$ is the Bohr radius, and $n_{i}$ and $n_{f}$ are the prinicipal quantum numbers corresponding to the initial and final states, respectively. For absorption, $n_{f}>n_{i}$ and $\Delta E>0$; however, for emission, $n_{f}<n_{i}$ and $\Delta E<0$. Taking $|\Delta E|=\hbar\omega_{o}$, the parameters $\omega_{o}$ and $\mathbf{k_{o}}$} are determined from Eq. (6) that gives
$\omega_{o}=\frac{R_{E}}{{n_{i}^{2}n_{f}^{2}}\hbar}|n_{f}^{2}-n_{i}^{2}|\ ,$ (7)
@@
${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{{n_{i}^{2}n_{f}^{2}}c\hbar}|n_{f}^{2}-n_{i}^{2}|{\mathbf{\hat{k}}_{o}}\ ,$ (8)
which guarantee that $\omega_{o}$ and $\mathbf{k_{o}}$ are positive for both absorption and emission.
Then, Eq. (5) becomes
@@ 3.3 Temporal and spatial solutions
To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$. Then, Eq. (9) becomes
@@
where $n$ is the principal quantum number; it is seen that the time-independent NAE directly displays Bohr’s rule. Since the solutions of both equations depend on $n$, let $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$. Note that $n=n_{f}$ for absorption, and $n=n_{i}$ for emission.
@@
[…]
and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 4.
Before solving Eq. (11), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$ for the absorbed EM radiation, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$. Similarly, for EM radiation emitted by an atom, the direction of such radiation is undetermined until it is measured by macroscopic instruments; therefore, $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$ remains also valid for emission.
It is also convenient to transform Eq. (11) to the new independent variable $r_{a}=r/a_{B}$, which gives
@@
According to Eq. (15), $\eta_{n}(r)$ reaches its maximum value when $r_{a}=n^{2}$ or $r=n^{2}a_{B}$, which is Bohr’s rule for quantum jumps (e.g., [3-5]). In other words, Eq. (15) is the mathematical form of Bohr’s rule, and it gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
Having obtained the time-independent NAE given by Eq. (15), it is seen that its form is different than a typical time-independent SE used in applications in NRQM (e.g., [4,5]). By being a second-order ODE, the SE can be considered as a Sturm-Liouville problem with its resulting eigenvalue equation; however, the NAE being a first-order ODE cannot be cast in the same form. Instead, the NAE requires specifying its eigenvalues $\omega_{o}$ and $\mathbf{{k}_{o}}$ as described above. Moreover, Eq. (15) directly displays Bohr’s rule, and its shows that the the time-independent wavefunction $\eta_{n}(r)$ reaches its minimum at $r=n^{2}a_{B}$. Bohr’s rule can also be obtained from the time-independent SE through the construction of hydrogen wave functions that satisfy the following condition $1/r>=1/(n^{2}a_{B})$ (e.g., [4,5]).
@@
$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{B}}\ \exp{\left[-\left(\frac{\beta_{B}}{n^{2}}\right)r_{a}\right]}\ ,$ (17)
@@
[…]